OpenStudy (anonymous):

Just want to be sure I am doing this right... ln4x=6...

OpenStudy (anonymous):

\[4^6=x\]

OpenStudy (anonymous):

Or x=16384

OpenStudy (anonymous):

That seems way too large to be right.

OpenStudy (anonymous):

It's incorrect x = 4096

OpenStudy (anonymous):

but ln 4x = 6 is 4x = e^6 x = e^6 / 4

OpenStudy (anonymous):

so its not 4^6

OpenStudy (anonymous):

Wait, how so?

OpenStudy (anonymous):

I agree with x = e^6 / 4, but thenwhy does x = 4096?

OpenStudy (anonymous):

it's not. 4^6 = 4096.

OpenStudy (anonymous):

But aren't you supposed to get rid of a log in order to solve for x?

OpenStudy (anonymous):

lnx means log base e of x so ln4x = log(base e)4x

OpenStudy (anonymous):

I know ln has a base of e, and that would be \[e^4x=6\]

OpenStudy (anonymous):

it messed up, one sec

OpenStudy (anonymous):

nvm, it won't let me do it. e^(4x)=6 is what I was trying to say

OpenStudy (anonymous):

Eh, I still don't get this. Could you rewrite this as \[e^6=4x\]

OpenStudy (anonymous):

divide both sides by 4, and you get: \[e^6/4=x\]

OpenStudy (anonymous):

which is exactly what you said...